What is the mass of gas at a pressure of 500 kPa and a temperature of 50 °C if it occupies a volume of 0.10 m3? The gas has a characteristic gas constant of 189 J kg-1 K-1
1
Expert's answer
2020-12-15T07:51:16-0500

A characteristic gas constant is the universal gas constant divided by the molar mass of the gas:

"R_c = \frac{R}{M}" .

The ideal gas law links the number of the moles of the gas , its volume, temperature and pressure:

"pV = nRT" .

Finally, the mass of gas is its number of the moles times molar mass:

"m = n\cdot M" .

Using these equations, the mass of the gas is:

"m = n\cdot M = \frac{pV}{RT}\cdot M = \frac{pV}{R_cT}"

"m = \frac{500\cdot10^3\cdot 0.10}{189 \cdot(50+273.15)} = 0.819" kg.


Remark: the units in the equation cancel and give kg as the result.


Answer: the mass of the gas is 0.819 kg.